OpenStudy (anonymous):

The temperature of a chemical reaction ranges between 30°C and 70°C. The temperature is at its lowest point when t = 0, the temperature is 90 degrees and increasing. What is a sine function that would model this reaction

OpenStudy (anonymous):

possible answers: f(x) = 20 sin 10x + 50 f(x) = 50 sin 10x + 20 f(x) = 50 sin pi over 5x + 20 f(x) = 20 sin pi over 5x + 50

OpenStudy (anonymous):

y = sina(bx+c)+k

OpenStudy (anonymous):

I already know that. How do you apply it to this problem? I think the question is bad and there is missing information